OpenStudy (anonymous):

Convert the following repeating decimal to a/b form where a and b are integers and b does not = 0 4.344444......

myininaya (myininaya):

Example: 5.35555555555555................ =5+.355555555555555555555555555 Let x=.355555555........................ Now do 10x which is 10(.355555555555555...............) I chose to multiply by 10 because after that that is where the repeating began 10x=3.5555555555555555555555555555 And do 10x-x=3.5555555555555555555555555....-.35555555555555555555.... 9x=3.55-.35 9x=3.2 x=3.2/9 But you want an integer over an integer So you need to multiply top and bottom by 10 like this: x=3.2(10)/[9(10)] x=32/90 Now recall we have 5.355555555555555............ =5+.35555555555555555555555 So we have =5+32/90 You can combine the fractions by finding a common denominator

myininaya (myininaya):

Let's look at yours when you have read through that example.